OpenStudy (anonymous):

Fill in the blank with a digit such that the resulting number is divisible by 11. (a) 962,974,__34 (b) 92,818,__21 (c) 57,13__,473

OpenStudy (maddy1251):

Does it mean the number following will be able to be divided by 11?

OpenStudy (markchernioglo):

what do you suppose the answer is?

OpenStudy (maddy1251):

@markchernioglo I will watch you do this, because I think I know the answer, but not too sure.. :p

OpenStudy (markchernioglo):

lol same but ill try

OpenStudy (maddy1251):

I believe I know what it is..

OpenStudy (markchernioglo):

so i think A. could be eliminated becuase 962,974 is not a factor of 11.

OpenStudy (maddy1251):

(a) 962,974,__34 When divided by 11, we don't get a whole number. (b) 92,818,__21 When divided by 11, we don't get a whole number. (c) 57,13__,473 When divided by 11, we get 43.

OpenStudy (markchernioglo):

yup maddie you got it.

OpenStudy (markchernioglo):

that is correct

OpenStudy (anonymous):

for the question its asking to find the missing blank number

OpenStudy (maddy1251):

@markchernioglo I took the thunder, you were starting tho.

OpenStudy (markchernioglo):

:)

OpenStudy (anonymous):

to make it true for it to be divided by 11

OpenStudy (maddy1251):

It is the last one, the number must be 43. Right @markchernioglo

OpenStudy (markchernioglo):

yup

OpenStudy (markchernioglo):

correct

OpenStudy (maddy1251):

Thought so. :-)

OpenStudy (maddy1251):

C is your correct answer @Jackiemelbs

OpenStudy (markchernioglo):

yup

OpenStudy (maddy1251):

If you want to go the extra mile, the answer you get from dividing that by 11 should be the fill in the blank spot.

OpenStudy (markchernioglo):

well igtg if anyone needs help maddy you are smart :)

OpenStudy (anonymous):

its not asking for which one is correct, for each letter it wants you to put the 1 missing number to make it true that the number can be divided by 11

OpenStudy (maddy1251):

@markchernioglo trust me you are uber smart too!! You were right there with me :-) I will fan you!

OpenStudy (markchernioglo):

lol same

OpenStudy (maddy1251):

So whatever you get from dividing by 11, they want in there?

OpenStudy (maddy1251):

Or that the answer is divisable by 11?

OpenStudy (anonymous):

no its something with some sort of divisibility rule of 11 that we have to use in order to find it and thats why i posted cause i get stuck on it

OpenStudy (maddy1251):

Ah okay. One second (-: Sorry, I misinterpretted.

OpenStudy (maddy1251):

Does it request any special numbers, is there a rule to these numbers in the sequence?

OpenStudy (anonymous):

Here an easy way to test for divisibility by 11. Take the alternating sum of the digits in the number, read from left to right. If that is divisible by 11, so is the original number. So, for instance, 2728 has alternating sum of digits 2-7+2-8 = -11. (I've tried using this but I can never get it to work)

OpenStudy (anonymous):

i guess what i don't understand is how to determine this if you are missing a number

OpenStudy (maddy1251):

Well, I understand that. But, you could honestly put in any numbers. There must be a catch,

OpenStudy (maddy1251):

Fill in the blank with a digit such that the resulting number is divisible by 11 (a) 962,974,__34 (b) 92,818,__21 (c) 57,13__,473 What I think is they want us to add up all the numbers up until the point of the blank spot possibly

OpenStudy (maddy1251):

Like for C, 5+7+1+3=16, but see that is not divisible by 11..

OpenStudy (maddy1251):

and 57 + 13 = 70 which is not divisible by 11, either.

OpenStudy (anonymous):

right....ugh this problem sucks haha

OpenStudy (maddy1251):

It does! I mean, I understand the rule. But there is a trick within these numbers.

OpenStudy (maddy1251):

And there is no pattern, the numbers are inconsistent.

OpenStudy (anonymous):

right....oh well i guess I'll just ask my professor cause i honestly have no idea

OpenStudy (maddy1251):

Its gotta do with the numbers, if they were in sequence, it would seem more reasonable, but the numbers fluctuate a lot. So, that doesn't help. you could pick any number under these circumstances, for it to be divisble by '11', unless there was something in your lesson that went over these.

OpenStudy (maddy1251):

Well I have one proposition

OpenStudy (maddy1251):

We could take each end number, such as a) for 34, times it by 11, giving us 374, putting that into the blank spot.

OpenStudy (maddy1251):

That is a multiply of 11, and dividing it by 11 yields 34.

OpenStudy (maddy1251):

b) 21 x 11 = 231. 231 is a factor of 11, so it could be a possible answer.

OpenStudy (maddy1251):

And mybe put in 43 for c, because that multiplies out to give us 473.

OpenStudy (maddy1251):

So, A) 374 B) 231 c)43 ( maybe )

imqwerty (imqwerty):

do u know the divisibility test for 11

OpenStudy (maddy1251):

@imqwerty am I somewhat right? Haha

imqwerty (imqwerty):

wait a sec ima see :) i ddin't read anything i jst cmnted cx

OpenStudy (maddy1251):

Haha okay XD

imqwerty (imqwerty):

what digits are u filling in the 1st one? i dun get it u wrote 374 but there are only 2 spaces

OpenStudy (maddy1251):

A) 374 B) 231 c)43 Are my answers

OpenStudy (maddy1251):

I disregarded the spaces, I think it was just spaces to be space

OpenStudy (maddy1251):

I highly doubt c is 5203. That seems to be too large.

OpenStudy (maddy1251):

However, if you divide 473 x 11 = 5203. Which then divided by 11 yields 473, so, that is a possibility.

OpenStudy (maddy1251):

multiply, not divide..

OpenStudy (maddy1251):

A) 374 B) 231 c)43 or 5203.

OpenStudy (anonymous):

a) 0 b) 1 c) 4 Works: a) 9 -6 + 2 - 9 + 7- 4 + x - 3 + 4 = closest number divisible by 11 0 + x = closest number divisible by 11, in this case it is 0 0 + x = 0, x = 0 b) 9 - 2 + 8 - 1 + 8 - x +2 -1 = closest number divisible by 11 23 - x = closest number divisible by 11, in this case it is 22 23 - x = 22, x = 1 c) 5 - 7 + 1 - 3 + x - 4 + 7 - 3 = closest number divisible by 11 -4 + x = closest number divisible by 11, in this case it's 0 -4 + x = 0, x = 4

OpenStudy (anonymous):

(a) 962,974,034 (b) 92,818,121 (c) 57,134,473

OpenStudy (anonymous):

Refer to the attached Mathematica calculation.